# Organizing Unstructured Image Collections using Natural Language

### v1

@@ 5 Applications
Discovering and mitigating dataset bias: Given an image collection that contains spurious correlations (Geirhos et al., 2020), we are curious whether we can proactively find this issue caused by data bias directly from the training images without relying on either the annotations (Sagawa et al., 2020) or post hoc misclassified images (Kim et al., 2024).
As a case study, we applied the proposed TeDeSC framework to the 162k training images of the CelebA (Liu et al., 2015) dataset—a binary hair color classification dataset where the target label “Blond” is spuriously correlated with the demographic attribute “Female” in its training split. Additional details are provided in App. I.1.
Findings: As expected, our method identified the grouping criteria Hair color and Gender. Next, we analyzed the predicted gender distributions within the “Blond” and “Not Blond” (all other colors) clusters. As shown in Fig. 7(a), we observed that the gender distribution within the “Blond” cluster is highly skewed, with 86.5% of the images representing females, closely matching the ground-truth distribution (94.3%). Such an imbalance confirms the potential issue of spurious correlations between “Blond” and “Female”. To further validate this observation, following B2T (Kim et al., 2024), we used our predicted distributions to train a debiased model with GroupDRO (Sagawa et al., 2020) and compared it with other unsupervised bias mitigation methods, including JTT (Liu et al., 2021), CNC (Zhang et al., 2022), B2T, and GroupDRO trained with ground-truth labels. As shown in Fig. 7(b), our debiased model achieved robust performance, comparable to that of B2T, demonstrating the reliability of its discovered distributions.
Figure 8: Bias discovery results in T2I models. Bias evaluation results are shown for the associated occupations.
@@
Discovering novel bias in Text-to-Image Diffusion Models: Stereotypical biases related to gender or race (Naik & Nushi, 2023) in images generated by widely accessible text-to-image (T2I) models like Stable Diffusion XL (SDXL) (Podell et al., 2024) and DALL·E3 (Betker et al., 2023) have sparked intense discussions (Nicoletti & Bass, 2023) and significant research (Bianchi et al., 2023). However, beyond these well-known biases extensively studied in social science, we wonder: Are there other biases present in T2I-generated images? To explore this question, we selected nine occupations (e.g.,, Nurse or CEO) from prior studies (Bianchi et al., 2023; Bolukbasi et al., 2016) as subjects for our case study on bias. For each occupation, we generated 100 images using the prompt “A portrait photo of a $<$OCCUPATION$>$” with both DALL-E3 and SDXL, resulting in a total of 1.8k images (see Fig. 20 and Fig. 21 for examples). We applied TeDeSC to these images and automatically discovered 10 grouping criteria (possible bias dimensions) along with their predicted distributions for each occupation image set. To inspect the biases, we calculated the normalized entropy of each distribution to quantify the bias intensity, as proposed by D’Incà et al. (2024), and identified the dominant cluster—the one containing the most images—as the possible bias direction. Furthermore, we conducted a user study with 54 anonymous participants to assess how well our bias discovery results align with human judgment. We found that our system’s bias intensity scores closely matched human ratings—with an Absolute Mean Error (AME) of 0.1396 on a scale from 0 to 1—and that its predicted bias directions agreed with human evaluations 72.3% of the time. We highlight some key findings below and provide complete results and evaluations in App. I.2.
Findings: In Fig. 8, we present some of our notable findings. Without any predefined biases, our method successfully identifies well-known social biases associated with certain occupations. As shown in Fig. 8(a–c), the SDXL-generated images exhibit near-total social bias for occupations like Nurse, Firefighter, and Basketball Player, displaying gender and racial imbalances that exceed official statistics such as those from U.S. Bureau of Labor Statistics (2021). In contrast, DALL·E3 shows markedly improved performance in mitigating social biases—likely due to its bias mitigation “guardrails” (OpenAI, 2022). More importantly, as shown in Fig. 8(d–f), our method revealed previously unnoticed novel bias dimensions where both models exhibit moderate biases. For example, we found that SDXL strongly associates CEOs with “Grey” hair color, while DALL·E3 favors “Dark” hair. Interestingly, and contrary to trends observed in well-known social biases, DALL·E3 shows stronger biases than SDXL in Hair style and Grooming for Nurse and Teacher. For instance, 91% of the Teacher images generated by DALL·E3 depict individuals who are clean-shaven (Fig. 8(f)). The gap in DALL·E3’s performance between addressing well-known and novel biases suggests that even industrial T2I models with guardrail systems may overlook some novel biases, underscoring the importance of thoroughly examining biases beyond the commonly recognized ones.
Figure 9: Analysis of social media photo popularity on SPID dataset. We show the viral and major popular clusters along with the popularity distribution of data points within these clusters across five criteria (in Grey).
Analyzing Social media image popularity: Given a set of social media photos with significantly varying popularity for each photo, it is natural to ask: What visual element makes a photo popular? As a case study, we applied TeDeSC to 4.1k Flickr photos from the SPID dataset (Ortis et al., 2019), each accompanied by a popularity score based on the number of views. After our method grouped the photos into semantic clusters based on 10 discovered criteria, we calculated two scores for each cluster: i) $\bar{P}$: the average popularity of photos within the cluster, and ii) $\bar{P}_{w}$: the average popularity $\bar{P}$ weighted by the cluster’s proportion relative to the entire dataset. Here, we consider the cluster with the highest $\bar{P}$ as viral, meaning its photos are highly popular but may be few in number, and the cluster with the highest $\bar{P}_{w}$ as major popular, where the majority of popular photos are concentrated. We discuss the key findings below and provide complete results in App. I.3.
Findings: As shown in our analysis in Fig. 9, by combining the interpretable groupings from our method with popularity scores, the visual elements that drive virality (clusters with the highest $\bar{P}$) and the common traits shared among widely popular images (clusters with the highest $\bar{P}_{w}$) across different dimensions clearly emerged. Interestingly, we further observe that viral visual elements often exhibit sharp semantic contrast with those of the majority of popular images, such as Musical activities vs. Rest and relaxation, and High-intensity expressions vs. Neutral emotion. This suggests that viral content is more attention-grabbing, likely due to novel, unexpected, or intense visuals that stand out, especially in an era where the average attention span is less than 9 seconds (McSpadden, 2015; Farid, 2024). During our analysis, as shown in Fig. 9, we unexpectedly found that some of the most popular images in certain clusters contained not safe for work (NSFW) content, which had not been previously identified in the SPID dataset. This suggests that some images gain popularity due to negative or provocative visuals. It also highlights the need for more thorough inspection of visual datasets to avoid harm, and our proposed framework can be a useful tool for this purpose.
@@ 6 Conclusion
In this work, we introduce the task of semantic multiple clustering and propose TeDeSC, a system that automatically discovers grouping criteria in natural language from large image collections and uncovers interpretable data substructures based on these criteria. We rigorously evaluate various design choices of TeDeSC on four existing and two newly proposed benchmarks, and demonstrate its ability to reveal valuable insights that might otherwise go unnoticed in various real-world applications.
@@ Acknowledgments
This work was supported by the MUR PNRR project FAIR - Future AI Research (PE00000013), funded by NextGenerationEU. E.R. acknowledges partial support from the PRECRISIS project, funded by the EU Internal Security Fund (ISFP-2022-TFI-AG-PROTECT-02-101100539), the EU project SPRING (No. 871245), and the PRIN project LEGO-AI (Prot. 2020TA3K9N). This work was conducted within the Vision and Learning joint laboratory of FBK and UNITN. We extend our gratitude to Feng Xue for his valuable suggestions on plot creation. M.L. would like to thank Margherita Potrich for her unwavering support.
@@
We do not anticipate any immediate negative societal impacts from our work. However, we encourage future researchers building on this work to remain vigilant, as we have, about the potential for TeDeSC, which integrates LLMs and MLLMs–particularly their human-like reasoning abilities– to be used both for good and for harm.
The motivation behind our studies on biases in existing datasets and text-to-image (T2I) generative model outputs, as well as our exploration of social media image popularity, is to reveal and address these biases and the presence of sensitive or not-safe-for-work (NSFW) content that objectively exist in the datasets and models. We emphasize that our aim is to study and mitigate these issues, and in doing so, we do not create any new biases or disturbing content. Specifically, in Sec. 5, we use well-established benchmarks, such as CelebA (Liu et al., 2015), for our study of dataset bias, and for bias discovery in T2I generative models, we select occupation-related subjects known to be associated with biases from prior studies (Bianchi et al., 2023; Bolukbasi et al., 2016). Furthermore, our framework reveals previously undisclosed sensitive and sexual content in the SPID dataset (Ortis et al., 2019). We responsibly present these findings in Sec. 5, applying significant blurring to disturbing content, with the intention of raising community awareness about the need to further scrutinize NSFW content in existing benchmarks. However, we acknowledge that our methodology and findings could potentially be misused by malicious actors to promote harmful narratives or discrimination against certain groups. We strongly oppose any such misuse or misrepresentation of our work. Our research is conducted with the aim of advancing technology while prioritizing public welfare and well-being.
@@
Lastly, we emphasize that our proposed framework, TeDeSC, relies on open-source LLMs and MLLMs, allowing full deployment on local machines. We refrain from using APIs from industrial LLMs or MLLMs, both to ensure reproducibility and to protect data privacy.
Reproducibility Statement
We will release all essential resources required to reproduce the experimental results presented in this project, including source code, exact prompts, benchmarks with their data splits, and generated images, upon publication. Our proposed framework, TeDeSC, is built on open-source, publicly accessible models to ensure reproducibility. In Sec. 3, we provide a detailed description of how our framework is constructed. Additionally, App. C contains further implementation details, including exact prompts, to help practitioners easily reproduce our method. Details regarding the implementation of the compared methods are also provided in App. D. Moreover, we present extended numerical experimental results in App. E, alongside comprehensive findings for the application study in App. I. We believe that the thorough descriptions of our methodology, the extensive presentation of experimental results, and the open-source nature of our framework ensure that this work is highly reproducible, enabling future researchers and practitioners to readily apply our method to various domains.
@@
This appendix provides detailed supplementary information supporting the implementations, experiments, and findings presented in the paper. First, in App. A, we offer an extensive discussion of Related Work, covering tasks and methods pertinent to our study. In App. B, we describe the benchmarks used in our study, including the construction of the newly proposed COCO-4c and Food-4c datasets, along with the process for creating hard ground-truth criteria for evaluating different proposers. App. C outlines the prompts and implementation details for our proposed framework, including both Criteria Proposers and Semantic Groupers. App. D provides a deep dive into the implementation specifics of the compared methods used in our experiments. In App. E, we present additional quantitative results that supplement the findings reported in the main paper, covering evaluations of the Criteria Proposer, Semantic Grouper, and comparisons with other clustering methods. App. F extends our analysis of the proposed framework, including system sensitivity and studies on fine-grained image collections. App. G presents further qualitative results for the predicted clusters, while App. H offers a more detailed analysis of the failure cases encountered by our method. App. I provides additional findings, along with implementation details and user study results for the three applications explored in this work. Finally, App. J presents a deeper discussion of related settings pertinent to our research.
We will release all essential resources for reproducing this work, including code, prompts, benchmarks, and annotations, upon publication.
Appendix A Related Work
Image Clustering. Image clustering (Xu & Wunsch, 2005) discovers hidden grouping structures within large, unstructured, and unlabeled image collections, serving as a tool for various data-driven applications (Wazarkar & Keshavamurthy, 2018). To achieve this, deep clustering (DC) methods such as DEC (Xie et al., 2016) and SCAN (Van Gansbeke et al., 2020) focus on simultaneously learning feature representations and cluster assignments using deep neural networks via self-supervised techniques (Caron et al., 2020; Zhong et al., 2021; Ren et al., 2024). Furthermore, large-scale pre-trained feature representations like DINOv1 (Caron et al., 2021), DINOv2 (Oquab et al., 2023), and CLIP (Radford et al., 2021) have also been shown to be effective at clustering image collections in a zero-shot fashion with the help of KMeans (Vaze et al., 2022; Liu et al., 2024f; Han et al., 2023).
However, it is well-known that “clusters are in the eye of the beholder” (Estivill-Castro, 2002); there often exist multiple ways to partition the same image collection into clusters, and what constitutes a cluster depends on the user’s needs. This insight has led to the study of Multiple Clustering (MC) (Ren et al., 2022; Yao et al., 2023; Yu et al., 2024), which aims to simultaneously learn feature representations and cluster assignments from different perspectives to find various ways of grouping the same data, enabling alternative interpretations from different viewpoints. Although existing MC methods have achieved impressive results (Yao et al., 2023) on some benchmarks, they share similar limitations with deep clustering approaches (Xie et al., 2016; Van Gansbeke et al., 2020). Their results require intensive manual post-analysis, and they also hold strong assumptions: users must specify (i) the number of potential clusterings and (ii) the number of clusters within each clustering. However, when dealing with millions of unstructured images, it is infeasible for users—who are trying to understand the data—to know this information a priori.
@@ Appendix A Related Work
In stark contrast to prior work, we introduce and study the task of semantic multiple clustering (SMC). Instead of requiring users to specify the grouping criteria, SMC seeks to actively and automatically discover criteria expressed in natural language from large visual data and uncover the corresponding semantic substructures, without access to any of the aforementioned human priors. As demonstrated in Sec. 5, the flexibility provided by our proposed SMC framework adds significant value to various data-driven applications, unveiling novel insights about the data that might not have been noticed before.
Topic Discovery. The setting of semantic multiple clustering (SMC) is also related to the field of Topic Discovery (Blei et al., 2003; Wang et al., 2009; Eklund & Forsman, 2022) in natural language processing, which aims to identify textual themes from large text corpora (e.g.,, documents). Our work shares motivational similarities with topic discovery because both tasks seek to find common, thematic concepts from large volumes of data. In contrast, our work focuses on discovering thematic criteria from large visual content. However, indeed, the core challenges of SMC and topic discovery are highly similar: they both require systems that can concurrently reason over large volumes of data. Nevertheless, SMC is an even more challenging task than topic discovery for two reasons: i) semantics are not explicitly expressed in images, whereas they are in text; ii) there is currently no vision model that can encode large sets of images and reliably reason over them. Thus, in this work, we translate images to text and use text as a proxy to elicit the large-scale reasoning capability of large language models (Meta, 2024b).
Multimodal Large Language Model. Recent advancements in multimodal large language models (MLLMs) have been driven by the availability of large-scale vision-language aligned training data. The typical paradigm (Liu et al., 2024c) involves using a pre-trained large language model (LLM) (Meta, 2024a; Chiang et al., 2023; Jiang et al., 2023; Meta, 2024b) alongside a pre-trained vision encoder (Radford et al., 2021). A projector is learned to align visual inputs with the LLM in the embedding space, which enhances visual understanding by utilizing the reasoning capabilities of LLMs. Several models have achieved significant success in zero-shot image captioning and visual question answering (VQA), including BLIP-2 (Li et al., 2023), BLIP-3 (Xue et al., 2024), Kosmos-2 (Peng et al., 2023), and the LLaVA series (Liu et al., 2024c; b; Li et al., 2024). In our proposed TeDeSC framework, we employ MLLM primarily as a zero-shot image parser, converting visual information into text and using this text as a proxy to elicit LLMs for reasoning over large image collections and discovering grouping criteria. Additionally, we leverage the multi-image reasoning capability of LLaVA-NeXT-Interleave (Li et al., 2024) to establish a baseline image-based proposer for the SMC task, while utilizing BLIP-2 with customized prompts in a VQA style (Shao et al., 2023; Zhu et al., 2023) as the image-based grouper to form semantic clusters linked to specific visual content within the images.
[…]
Text-Driven Image Retrieval. Given a query text (e.g.,, “sofa” or “person wearing a blue T-shirt”), text-driven image retrieval methods (Karthik et al., 2024; Liu et al., 2023; Wu et al., 2023) aim to find images from an image collection that are relevant to the query. In other words, in the scenario we are considering, given the image collection and a list of text queries, one can organize images according to the text using text-driven image retrieval techniques. In this context, the query can be considered as a sort of “cluster name”. However, this differs significantly from the proposed task of semantic multiple clustering (SMC), because SMC requires both discovering the textual criteria and the corresponding textual clusters. Thus, without knowing text queries as prior information, text-driven image retrieval methods are not able to accomplish SMC.
Appendix B Benchmark Details
B.1 Benchmark Construction of COCO-4c and Food-4c

### v4

@@ 6 Applications
In this section, we apply the best caption-based $\mathcal{X}$-Cluster on three applications to showcase its ability to surface novel, human-interpretable insights.
@@
Discovering Novel Bias in Text-to-Image (T2I) Diffusion Models: Do T2I models encode biases beyond the well-studied gender and race stereotypes? (Naik and Nushi, 2023) To investigate this, we selected nine occupations (e.g., Nurse, CEO) from prior studies (Bianchi et al., 2023; Bolukbasi et al., 2016) and generated 100 images per occupation via the prompt “A portrait photo of a <OCCUPATION>” using DALL-E3 (Betker et al., 2023) and SDXL (Podell et al., 2024), resulting in 1.8k images. Applying $\mathcal{X}$-Cluster, we automatically identified 10 grouping criteria (bias dimensions) and their distributions for each occupation. To quantify bias, we measured the normalized entropy of each dimension’s distribution (D’Incà et al., 2024) as bias intensity and marked the dominant (largest) cluster as the potential bias direction. We conducted a user study with 54 participants to validate our results: $\mathcal{X}$-Cluster’s predicted bias intensity closely matched ratings from the interviewees with an Absolute Mean Error of 0.1396 on a 0–1 scale, and bias directions agree 72.3% of the time with the interviewees. See App. P.1 for user study details.
Findings: As shown in Figure 8, our method identifies both well-known and previously unnoticed novel biases without relying on predefined categories. For SDXL, roles such as Nurse, Firefighter and Basketball Player show pronounced gender and racial skews that exceed labor-force statistics (U.S. Bureau of Labor Statistics, 2021), whereas DALL-E3’s guardrails (OpenAI, 2022b) mitigate many of these effects. More notably, Figure 8(d–f) highlights previously unnoticed bias dimensions. For example, SDXL strongly associates CEOs with “Grey” hair, while DALL·E3 favors “Dark” hair. Additionally, DALL·E3 shows stronger biases in Hair style and Grooming for occupations like Nurse (Figure 8(e)) and Teacher (Figure 8(f)). These findings suggest that while industrial T2I models with guardrails may have mitigated well-known biases, they often overlook emerging or less-discussed ones – highlighting the need for broader audits, where $\mathcal{X}$-Cluster can serve as a valuable tool. See further findings in App. P.1.
Analyzing Social Media Image Popularity: What visual traits drive popularity? To explore this, we apply $\mathcal{X}$-Cluster to 4.1k Flickr photos from the SPID dataset (Ortis et al., 2019), where popularity is measured by image view count. $\mathcal{X}$-Cluster discovered ten grouping criteria and organizes photos into semantic clusters under each. Using the grouping results, Figure 9 compares the sample popularity distributions of the Top Trending and the Top Mainstream clusters for three criteria.
Findings: As shown in Figure 9, combining $\mathcal{X}$-Cluster’s grouping with popularity scores provides a direct interpretation of the visual elements that drive trends versus those that define widely uploaded images. Interestingly, we find that trending elements often contrast with mainstream ones, such as “Musical activities” vs. “Rest and relaxation” or “High-intensity expressions” vs. “Neutral emotion”. These results suggest that attention-grabbing visuals stand out due to novelty or intensity, especially in today’s short attention span era (McSpadden, 2015; Farid, 2024), underscoring $\mathcal{X}$-Cluster as a powerful tool for deep dataset analysis and understanding social behavior. For full findings and additional analysis, see App. P.2.
Confirming and Mitigating Dataset Bias: Lastly, in App. P.3, we used $\mathcal{X}$-Cluster to confirm the gender bias in 162k CelebA (Liu et al., 2015) training images and successfully trained a debiased model using the predicted groupings with DRO (Sagawa et al., 2020). Tab. 35 shows our debiased model achieves robustness comparable to B2T (Kim et al., 2024), which uses annotations for bias detection.
Figure 9: Social media image popularity analysis. We show the popularity score distributions for Top Trending (have highest average popularity score) and Top Mainstream (contain most images) clusters, discovered by $\mathcal{X}$-Cluster across three criteria.
@@ 7 Conclusion
We introduced the OpenSMC task and proposed $\mathcal{X}$-Cluster for discovering grouping criteria and semantic substructures in an unstructured image collections, expressed in natural language. We evaluated $\mathcal{X}$-Cluster on six benchmarks and three applications, showing its ability to automatically discover valuable insights stored in visual data.
@@ Acknowledgments
This work was supported by the MUR PNRR project FAIR - Future AI Research (PE00000013), funded by NextGenerationEU. E.R. acknowledges partial support from the PRECRISIS project, funded by the EU Internal Security Fund (ISFP-2022-TFI-AG-PROTECT-02-101100539), the EU project SPRING (No. 871245), and the PRIN project LEGO-AI (Prot. 2020TA3K9N). This work was conducted within the Vision and Learning joint laboratory of FBK and UNITN. We thank OpenAI for supporting our experiments through the Researcher Access Program (0000004498). We extend our gratitude to Feng Xue for his valuable suggestions on plot creation. M.L. would like to thank Margherita Potrich for her unwavering support.
Limitations
As discussed in § 5.1 and § 5.2, $\mathcal{X}$-Cluster still has significant room for improvement. It is thus not guaranteed to uncover all valid grouping criteria or perfectly align semantic substructures with user preferences. Moreover, although designed as an automatic, assistive tool for large-scale visual data analysis, $\mathcal{X}$-Cluster requires human oversight; users bear the ultimate responsibility for correctly interpreting the generated results.
Since $\mathcal{X}$-Cluster relies on multimodal large language models (MLLMs) and large language models (LLMs), it inherits their limitations, including potential errors, biases, or hallucinations. Foundation models such as LLaVA-Next (Liu et al., 2024b), CLIP (Radford et al., 2021), and Llama series (Meta, 2024b, a) reflect biases present in their internet-scale training data (Goh et al., 2021; Schramowski et al., 2022; Kirk et al., 2021; Yamada et al., 2023; Wan et al., 2023). These biases can be challenging to identify or quantify due to the vast—and sometimes proprietary—nature of the training sets (Radford et al., 2021; OpenAI, 2022a, 2023).
Furthermore, both MLLMs and LLMs may suffer from hallucinations. For LLMs (Ji et al., 2023; Yao et al., 2023a), this means generating information that appears accurate but is not. For MLLMs (or VQA models) (Ge et al., 2024), it involves creating image descriptions that seem reasonable but do not accurately reflect the visual content.
To alleviate these limitations, we provide detailed analyses of failure cases (App. I) for our proposed $\mathcal{X}$-Cluster framework, strategies for handling hallucinated or invalid criteria (App. K), and methods to mitigate biases (App. L) in the technical appendix.
Lastly, due to the zero-shot nature of $\mathcal{X}$-Cluster, its limitations may be mitigated by incorporating improved foundation models (OpenAI, 2025), employing bias alleviation techniques (Kong et al., 2024), or applying hallucination reduction strategies (Sahoo et al., 2024; Ge et al., 2024).
@@
We do not anticipate any immediate negative societal impacts from our work. However, we encourage future researchers building on this work to remain vigilant, as we have, about the potential for $\mathcal{X}$-Cluster, which integrates LLMs and MLLMs–particularly their human-like reasoning abilities– to be used both for good and for harm.
The motivation behind our studies on biases in existing datasets and text-to-image (T2I) generative model outputs is to reveal and address these biases that objectively exist in the datasets and models. We emphasize that our aim is to study and mitigate these issues, and in doing so, we do not create any new biases or disturbing content. Specifically, in § 6, we use well-established benchmarks, such as CelebA (Liu et al., 2015), for our study of dataset bias, and for bias discovery in T2I generative models, we select occupation-related subjects known to be associated with biases from prior studies (Bianchi et al., 2023; Bolukbasi et al., 2016). However, we acknowledge that our methodology and findings could potentially be misused by malicious actors to promote harmful narratives or discrimination against certain groups. We strongly oppose any such misuse or misrepresentation of our work. Our research is conducted with the aim of advancing technology while prioritizing public welfare and well-being.
@@
Lastly, we emphasize that our proposed framework, $\mathcal{X}$-Cluster, relies on open-source LLMs and MLLMs, allowing full deployment on local machines. We refrain from using APIs from industrial LLMs or MLLMs, both to ensure reproducibility and to protect data privacy.
@@
Overview
This supplementary material provides additional details supporting the implementations, experiments, findings, and discussions in the main paper.
In App. A, we first provide a reproducibility statement to ensure the transparency and replicability of our work.
App. B expands on Related Work, covering relevant tasks and methods. App. C describes the benchmarks used in our study, including the construction of the newly proposed COCO-4c and Food-4c datasets, as well as the process for creating hard ground-truth criteria for proposer evaluation. App. D details the evaluation metrics used in this work.
App. E presents the prompts and implementation details for $\mathcal{X}$-Cluster, covering both Criteria Proposers and Semantic Groupers. App. F provides implementation specifics of the compared methods. Additional quantitative results, including evaluations of the Criteria Proposer, Semantic Grouper, and comparisons with other clustering methods, are in App. G.
For qualitative analysis, App. H presents further visualizations of predicted clusters, while App. I examines failure cases. App. J investigates the impact of multi-granularity clustering. App. K explores the effect of invalid (hallucinated) criteria on system performance, and App. L studies the influence of foundation model hallucinations and biases.
App. M analyzes the computational cost and runtime of $\mathcal{X}$-Cluster. App. N and App. O extend our analysis with studies on system sensitivity and fine-grained image collections. App. P provides additional findings, implementation details, and user study results for the three explored applications. App. Q discusses how LLMs enhance image clustering. Finally, App. R outlines potential future research directions for our proposed OpenSMC task.
The Table of Contents on the next page outlines the main topics in this supplementary material, with hyperlinks for direct navigation to each section.
Appendix A Reproducibility Statement
Upon publication, we will open-source all essential resources for reproducing this work. Specifically, we will provide the full code implementation of $\mathcal{X}$-Cluster, along with the exact prompts used in each module. Additionally, we will release our two newly proposed benchmarks, COCO-4c and Food-4c, including annotations for each grouping criterion. Lastly, we will provide the code for our evaluation protocol, experiments, and application studies.
Appendix B Additional Related Work
Topic Discovery. The setting of open-ended semantic multiple clustering (OpenSMC) is also related to the field of Topic Discovery (Blei et al., 2003; Wang et al., 2009; Eklund and Forsman, 2022) in natural language processing, which aims to identify textual themes from large text corpora (e.g., documents). Our work shares motivational similarities with topic discovery because both tasks seek to find common, thematic concepts from large volumes of data. In contrast, our work focuses on discovering thematic criteria from large visual content. However, indeed, the core challenges of OpenSMC and topic discovery are highly similar: they both require systems that can concurrently reason over large volumes of data. Nevertheless, OpenSMC is an even more challenging task than topic discovery for two reasons: i) semantics are not explicitly expressed in images, whereas they are in text; ii) there is currently no vision model that can encode large sets of images and reliably reason over them. Thus, in this work, we translate images to text and use text as a proxy to elicit the large-scale reasoning capability of large language models (Meta, 2024a).
Multimodal Large Language Model. Recent advancements in multimodal large language models (MLLMs) have been driven by the availability of large-scale vision-language aligned training data. The typical paradigm (Liu et al., 2024c) involves using a pre-trained large language model (LLM) (Meta, 2024b; Chiang et al., 2023; Jiang et al., 2023; Meta, 2024a) alongside a pre-trained vision encoder (Radford et al., 2021). A projector is learned to align visual inputs with the LLM in the embedding space, which enhances visual understanding by utilizing the reasoning capabilities of LLMs. Several models have achieved significant success in zero-shot image captioning and visual question answering (VQA), including BLIP-2 (Li et al., 2023a), BLIP-3 (Xue et al., 2024), Kosmos-2 (Peng et al., 2024), and the LLaVA series (Liu et al., 2024c, b; Li et al., 2025). In our proposed $\mathcal{X}$-Cluster framework, we employ MLLM primarily as a zero-shot image parser, converting visual information into text and using this text as a proxy to elicit LLMs for reasoning over large image collections and discovering grouping criteria. Additionally, we leverage the multi-image reasoning capability of LLaVA-NeXT-Interleave (Li et al., 2025) to establish a baseline image-based proposer for the OpenSMC task, while utilizing BLIP-2 with customized prompts in a VQA style (Shao et al., 2023; Zhu et al., 2024) as the image-based grouper to form semantic clusters linked to specific visual content within the images.
@@ Appendix B Additional Related Work
Text-Driven Image Retrieval. Given a query text (e.g., “sofa” or “person wearing a blue T-shirt”), text-driven image retrieval methods (Karthik et al., 2024; Liu et al., 2023; Wu et al., 2023) aim to find images from an image collection that are relevant to the query. In other words, in the scenario we are considering, given the image collection and a list of text queries, one can organize images according to the text using text-driven image retrieval techniques. In this context, the query can be considered as a sort of “cluster name”. However, this differs significantly from the proposed task of open-ended semantic multiple clustering (OpenSMC), because OpenSMC requires both discovering the textual criteria and the corresponding textual clusters. Thus, without knowing text queries as prior information, text-driven image retrieval methods are not able to accomplish OpenSMC.
Appendix C Benchmark Details
C.1 Construction of COCO-4c and Food-4c